Crisp Cookware's common stock is expected to pay a dividend of $2 a share at the end of this year (D1 = $2.00); its beta is 0.70. The risk-free rate is 4% and the market risk premium is 4%. The dividend is expected to grow at some constant rate gL, and the stock currently sells for $48 a share. Assuming the market is in equilibrium, what does the market believe will be the stock's price at the end of 3 years (i.e., what is )? Do not round intermediate steps. Round your answer to the nearest cent.
Answer #1

Required rate=risk free rate+Beta*market risk premium

=4+(4*0.7)=6.8%

Required return=(D1/Current price)+Growth rate

0.068=(2/48)+Growth rate

Growth rate=0.068-(2/48)

=0.0263333333

P3=Current price*(1+Growth rate)^3

=$48*(1+0.0263333333)^3

=$51.89(Approx).